### v1

@@ IV-B Dataset
Our primary dataset was sourced from Kaggle [50]. We performed pre-processing on corpus of approximately 112k articles for our experiments. The dataset has 4 unique categories i.e., Science and Technology, Business and Economics, Entertainment and Sports. Figure 5 displays a high-level overview of the articles split across different categories.
Figure 5: Graphical representation of categories by frequency of Urdu news dataset
@@ IV-C Dataset Preprocessing
The raw Urdu news data underwent a systematic pre-processing pipeline to ensure quality and consistency for subsequent analysis. This process consisted of five primary stages: (1) Dataset Loading and Encoding Normalization, where the initial corpus was loaded, and encoding issues were resolved to correct Urdu character representation. (2) Metadata Pruning and Composite Text Construction, involving the removal of non-informative attributes (e.g., author, date stamps not used in this study) and the concatenation of article headlines with corresponding body text to form a unified content field. (3) Text Cleaning and Normalization, which stripped HTML tags, punctuation, special characters, and extraneous whitespace to standardize the textual data. (4) Urdu Stop-Word Elimination, wherein a curated list of high-frequency, low-information Urdu words was filtered out to reduce noise. (5) Invalid and Redundant Record Filtering, where entries with null values, exact duplicates, or text length below a defined threshold were removed to finalize the dataset. Figure 6 illustrates provides a visual overview of the complete pre-processing pipeline.
Figure 6: Complete Urdu news dataset pre-processing pipeline
@@ IV-C1 Dataset loading and encoding correction
Urdu text datasets often suffer from character encoding inconsistencies when processed through standard file reading methods. This issue arises due to variations in file encoding formats and platform-specific text handling. To preserve the integrity of the Nastaliq script, which includes complex character combinations and diacritical marks, the dataset was initially loaded using the ISO-8859-1 encoding scheme.
A custom decoding function was implemented to detect and correct misencoded characters programmatically. This function operates by detecting encoding mismatches through character pattern analysis, re-encoding problematic text segments from ISO-8859-1 to UTF-8, and validating the corrected text against Urdu Unicode ranges.
This systematic correction ensured accurate representation of all Urdu characters, including special diacritics (such as zabar, zer, and pesh) and punctuation marks specific to Urdu script. The encoding normalization was applied uniformly across all text-bearing columns: Headline, News Text, Category, and Source. This foundational step was critical to prevent information loss during subsequent processing stages.
@@ IV-C2 Removal of irrelevant metadata and composite content creation
News datasets typically contain various metadata fields that, while potentially useful for some analysis, do not contribute to semantic understanding in text classification and embedding tasks. To reduce dimensionality and eliminate potential noise sources, non-informative metadata columns were systematically removed. These included: Index: Sequential identifiers with no semantic value, Date: Publication timestamps (not used in this text-focused study), URL: Web addresses containing platform-specific patterns, Source: Publication sources (retained only for provenance tracking), News Length: Derived statistics rather than raw content.
Following metadata pruning, a unified textual representation was created to capture comprehensive article semantics. The Headline and News Text columns were concatenated using a delimiter-aware joining method to form a new column named content. This composite field architecture preserves the concise, information-dense nature of headlines. It incorporates the detailed contextual information from article bodies and also creates a single input stream for embedding models, simplifying pipeline architecture.
The transformation ensures that models receive both the topic-signaling function of headlines and the explanatory depth of full articles. Figure 7 and Figure 8 illustrates the dataset structure before and after this transformation phase respectively.
@@ IV-C3 Text cleaning and normalization
Raw text from web sources often contains structural artifacts, formatting remnants, and inconsistent character usage that can adversely affect downstream processing. To create a clean, standardized corpus, a multi-layered text normalization pipeline was implemented:
a)- HTML and formatting tag removal: Web-scraped content frequently includes HTML tags, CSS classes, and JavaScript fragments. These were stripped using a combination of regular expressions and HTML parsing libraries, ensuring complete removal while preserving textual content.
b)- Non-Urdu character elimination: Given the focus on Urdu-language analysis, non-relevant character sets were removed: English alphabets (A-Z, a-z), arabic numerals (0-9), mixed-language characters and symbols, platform-specific artifacts (e.g., emoticons, emojis not relevant to news context).
c)- Punctuation normalization: Both Urdu-specific and generic punctuation marks were systematically removed to create word-level tokens without boundary interference: generic punctuation: ‘.‘ (period), ‘,‘ (comma), ‘;‘ (semicolon), ‘:‘ (colon), quotation marks, brackets of all types, Urdu comma, Urdu semicolon, Urdu question mark.
d)- Whitespace standardization: Multiple consecutive whitespace characters (spaces, tabs, newlines) were collapsed into single spaces, and leading/trailing whitespace was trimmed. This normalization ensures consistent token boundaries during subsequent processing.
The combined effect of these cleaning operations is a homogeneous Urdu text corpus where semantic content is preserved while structural noise is minimized. This clean text serves as optimal input for both traditional NLP pipelines and modern embedding models.
@@ IV-C4 Urdu stop-word elimination
Stop words (high-frequency function words that carry minimal semantic content) can dominate text representations without adding meaningful information. For Urdu, which has a rich set of grammatical particles and connectors, stop-word removal is particularly impactful. A manually curated list of 127 Urdu stop words was compiled through linguistic analysis and frequency distribution studies of the corpus. The stop-word list includes several categories of function words:
@@
The removal process was applied after text cleaning but before any tokenization or feature extraction. This ordering ensures that stop words are identified in their normalized form, increasing removal accuracy. The efficacy of the Urdu stop-word removal process is demonstrated in Figure 13, which presents three comparative examples. Each case shows the preprocessed text (before), the resulting text after stop-word filtering (after), and the consequent reduction in text length (in characters), highlighting the substantial compression achieved while preserving core semantic content.
Figure 13: Common adverbs and discourse markers in Urdu language
Quantitatively, stop-word removal reduced the corpus size by approximately 33.5%, indicating the high frequency of these function words in Urdu news text. This reduction decreases computational requirements while potentially improving model performance by focusing on content-bearing terms. Figure 14 shows all 127 stop words that were identified and removed from dataset.
@@ IV-C5 Removal of invalid and redundant entries
Data quality assurance is critical for building reliable machine learning models. The preprocessing pipeline includes systematic filtering to eliminate problematic records through three validation criteria:
a)- Null value detection and removal: Records with missing values in essential columns were identified and removed. The validation focused on: content column: Empty or null text fields, Category column: Unlabeled or ambiguous category assignments, and Critical metadata: Missing values in fields required for analysis. This step ensures that all training examples contain complete information for both features and labels.
b)- Duplicate content identification: Duplicate articles, whether from republication, syndication, or crawling artifacts, can bias model training by over-representing certain content. This process maintains dataset diversity while eliminating redundant information.
c)- Minimum length filtering: Extremely short text entries (fewer than 15 characters) typically lack sufficient semantic content for meaningful analysis. These may include broken or truncated articles, navigational or interface text mistakenly captured, and placeholder or template content. Such entries were removed to ensure all training examples contain substantive content.
@@ IV-C6 Post-Processing Dataset Structure
Following the complete pre-processing pipeline, the dataset is structured for optimal use in downstream tasks. The dataset is stored in a tabular format with UTF-8 encoding to ensure compatibility with Urdu text processing libraries and frameworks. This structured, cleaned dataset serves as the foundation for all subsequent tasks described in the following sections.
The combined effect of these filtering operations is a clean, consistent dataset optimized for machine learning applications. Table II summarizes the quantitative impact of the complete pre-processing pipeline on the dataset characteristics.
@@ IV-D Proposed Methodology
The ULTRA framework implements a semantic search-based recommendation system tailored for Urdu news. As formalized in Section IV-A, the core challenge is to retrieve semantically relevant articles from a corpus of over 112k Urdu news items based on user queries of varying lengths and semantic granularity. By ”query granularity,” we refer to the level of detail and specificity in a user’s search input—ranging from concise, keyword-like queries, to elaborate, descriptive sentences. However, in our framework, news articles or full-length documents predominantly function as long queries, targeted at the use case where recommendations are generated for content relevant to the article the user is currently viewing.
Our approach addresses this challenge through three interconnected components: (1) adaptive embedding generation that aligns with query granularity, (2) efficient dimensionality reduction for scalable retrieval, and (3) an optimized dual-path retrieval architecture. The complete methodology is designed to balance semantic accuracy with computational efficiency, making real-time recommendations feasible for large-scale deployment.
@@ IV-E Adaptive Embedding Generation
The foundation of semantic retrieval lies in generating high-quality vector representations that capture the meaning of both queries and documents. We employ a pre-trained Urdu transformer model, urduhack/roberta-urdu-small, which outputs 768-dimensional contextualized embeddings for input tokens [51]. To obtain a fixed-dimensional representation for an entire text sequence, pooling strategies are required to aggregate token-level embeddings.
@@ IV-E1 Pooling Strategy Analysis
Three common pooling techniques were implemented and evaluated: mean pooling, max pooling, and CLS pooling. Mean pooling computes the average of all token embeddings, effectively creating a centroid of semantic information. Max pooling extracts the maximum value across each dimension, potentially emphasizing salient features. CLS pooling utilizes the special classification token’s embedding, which is trained during pre-training to capture sentence-level semantics.
@@ IV-E2 Query-Adaptive Dual-Path Architecture
A key insight driving our methodology is the semantic granularity mismatch between user queries and article representations. Short, focused queries as in Figure 15.
Figure 15: Short length query example
typically express a specific topic or intent in few words, making them semantically more aligned with article headlines or titles of documents, which are themselves concise summaries, than with the distributed, detailed content of full articles. To address this, we implemented a query-adaptive routing mechanism based on character length, which dynamically selects the most appropriate semantic space for comparison.
@@
It is important to clarify that in the headline pathway, the system uses headline embeddings only for similarity matching. The actual recommendations presented to the user are the corresponding full articles. This design exploits the fact that headlines serve as accurate, dense semantic proxies for their articles, enabling efficient retrieval of relevant content without requiring comparisons against lengthy article bodies. The headline embedding for each article is stored alongside metadata containing its full content and category, ensuring that when a headline match occurs, the complete article is readily available for recommendation.
For content exceeding the model’s maximum sequence length of 512 tokens, a chunking strategy with 50-token overlap was implemented. The final content embedding was computed as the average of individual chunk embeddings, ensuring comprehensive semantic coverage of lengthy articles.
@@ IV-F Dimensionality Reduction for Scalable Retrieval
With the corpus scaling to around 112,000 articles, performing nearest-neighbor search in the original 768-dimensional space becomes computationally intensive. To enhance scalability while preserving semantic relationships, we investigated four dimensionality reduction techniques: Principal Component Analysis (PCA), Uniform Manifold Approximation and Projection (UMAP), Linear Discriminant Analysis (LDA), and Autoencoders.
@@ IV-F1 Evaluation Methodology
The performance of each reduction technique was assessed not in isolation, but by its fidelity to the original semantic space. For a given query, we compared the top-$k$ recommendations retrieved from the reduced-dimensional database against those from the original 768-dimensional database, which served as ground truth. The primary evaluation metric was the overlap score (the number of articles common to both result sets) providing a direct measure of how well the reduced space preserves retrieval rankings.
PCA, a linear technique maximizing variance preservation, consistently outperformed non-linear alternatives across both content and headline embeddings. Its deterministic nature and computational efficiency, combined with superior overlap scores, made it the preferred choice for both retrieval pathways.
@@ IV-F2 Dimensionality Optimization
Following the selection of PCA, we optimized the reduced dimension $d^{\prime}$ to balance retrieval accuracy with computational efficiency. For content embeddings (long query pathway), we evaluated dimensions of 64, 128, and 256 across queries of varying lengths (150–350 characters). The 128-dimensional reduction provided the optimal trade-off, maintaining high overlap with ground-truth recommendations while significantly reducing query latency.
For headline embeddings (short query pathway), evaluation at 100 character length of 100 queries revealed that 64 dimensions sufficed to capture the semantic essence of concise headlines without noticeable degradation in retrieval quality. This dimension-specific optimization reflects the differing information density between lengthy articles and succinct headlines.
The pooled and reduced embeddings of optimal dimensions are then stored in two Chroma vector DBs: vector DB containing content column embeddings and vector DB containing headline column embeddings.
@@ IV-G System Architecture and Implementation
The ULTRA framework implements the methodological components through an integrated pipeline illustrated in Figure 4. The system processes user queries through the following stages:
@@ IV-G1 Query Processing and Routing
Upon receiving a user query $q$, the system first computes its character length $\ell(q)$. Based on the threshold $\theta=150$ (remember this $\theta$ is specific to our recommender system based on news corpus, and may vary across different recommender systems), the query is routed to the appropriate pathway. Both pathways employ the same base transformer model but differ in their text source, pooling strategy, and reduced dimension.
@@ IV-G2 Embedding Generation and Retrieval
The system employs two separate vector databases, each tailored to a specific query type. The processing pipeline for a given query is determined solely by its character length $\ell(q)$ relative to the threshold $\theta=150$, which dictates the pooling strategy, dimensionality reduction configuration, and target database for similarity search.
For queries with $\ell(q)\geq\theta$ (long queries), the system follows the content pathway. The query text is encoded using mean pooling to generate a 768-dimensional embedding $\mathbf{v}_{q}$, which is then reduced to 128 dimensions via Principal Component Analysis (PCA). This reduced vector $\mathbf{v}^{\prime}_{q}$ is compared against a pre-computed database of content embeddings, each generated through the same mean pooling and PCA reduction process, using cosine similarity.
For queries with $\ell(q)<\theta$ (short queries), the system follows the headline pathway. The query undergoes CLS pooling to produce its initial embedding, which is then reduced to 64 dimensions via PCA. This compact representation is compared against a separate database of headline embeddings, each processed identically (CLS pooling + PCA to 64D), to find the most semantically similar headlines.
Both vector databases are implemented using ChromaDB with Hierarchical Navigable Small World (HNSW) indexing configured for cosine similarity, enabling efficient approximate nearest-neighbor search at scale.

### v2

@@ IV-B Dataset
The source of our dataset is Kaggle [50]. Our experiments used pre-processing of about a corpus of 112k articles. The data set includes 4 distinct categories i.e., Science & Technology, Business & Economics, Entertainment & Sports. An overview of the articles that are categorized into various categories in is presented in Figure 5.
Figure 5: Graphical representation of categories by frequency of Urdu news dataset [50]
@@ IV-C Dataset Preprocessing
The raw Urdu news data was subjected to a systematic pre-processing pipeline to guarantee quality and consistency to be used in further analysis. This was done through five major steps: (1) Dataset Loading and Encoding Normalization, during which the loading of the first corpus was conducted, and encoding problems were addressed to fix the representation of Urdu characters. (2) Metadata Pruning and Composite Text Construction, which entails eliminating non-informative metadata (e.g., author, date stamps not used in this study) and concatenation of article headlines with respective article text to create a single content field. (3) Text Cleaning and Normalization, which removed HTML tags, punctuations, special characters, and unnecessary whitespace to standardize the textual data. (4) Urdu Stop-Word Elimination, in which a list of high-frequency, low-information Urdu words was weeded to remove noise. (5) Invalid and Redundant Record Filtering, in which records that contained a null value, exact duplicate or those that had a text length less than a specified threshold were deleted to complete the dataset. Visual representation of the pre-processing steps are depicted in Figure 6.
Figure 6: Urdu news dataset pre-processing steps
@@ IV-C1 Dataset loading and encoding correction
The Urdu text datasets usually have a problem of inconsistencies in character encoding when handled using the normal file reading procedures. The problem is caused by the differences in file encoding formats and text processing that is platform-dependent. In order to maintain the integrity of the Nastaliq script with its complex combinations of characters and diacritical marks, the dataset was first loaded with the encoding scheme of the ISO-8859-1.
To identify and fix mis-encoded characters, a custom decoding function was used. This feature works by identifying the mismatch in the encoding based on the character pattern, re-encoding troublesome text snippets in the ISO-8859-1 to UTF-8, and comparing the corrected text with Urdu Unicode ranges.
This methodical correction guaranteed the correct representation of all Urdu characters, that include special diacritics (such as zabar, zer, and pesh) and punctuation signs characteristic of Urdu script. Encoding normalization was also done to all textual columns: Headline, News Text, Category, and Source. This was a crucial step to ensure that no information is lost in the later stages of processing.
@@ IV-C2 Removal of irrelevant metadata and composite content creation
News datasets are usually characterized by a range of metadata fields that, although they may be useful in certain analysis, do not help to understand the semantics in text classification and embedding tasks. Non-informative metadata columns were removed systematically to reduce the dimensions and possible noise factors. These were: Index: Sequential identifiers that are not semantically meaningful, Date: Publication dates (not used in this text-oriented study), URL: Web addresses with platform-specific patterns, Source: Publication sources (not used in this text-oriented study), and News Length: Derived statistics, not the raw content.
After metadata pruning, a single textual representation was developed to achieve in-depth article semantics. The Headline and News Text columns were joined together with a delimited joining technique to create a new column called content. This field is a composite which maintains the terse, information-rich headlines. It takes into account the contextual information of the article bodies in more detail, as well as builds a single input stream to embed models, which makes pipeline architecture simpler.
The change is such that the models are provided with topic-signaling role of headlines as well as the elaboration of full articles. Figure 7 and Figure 8 shows the structure of the dataset prior to this transformation step and after it respectively.
@@ IV-C3 Text cleaning and normalization
Web sources of raw text frequently include structural artifacts, remnant formatting and irregular use of characters which may have a negative impact on downstream processing. In order to produce a clean, standardized corpus, a multi-layered text normalization pipeline was used:
a)- HTML and formatting tag removal: The content scraped by the Web often contains HTML tags, CSS classes and JavaScript fragments. These were stripped with a mix of regular expression and HTML parsing libraries so as to remove them completely leaving behind textual content.
b)- Non-Urdu character elimination: Since it was about the analysis of Urdu characters, irrelevant sets of characters were eliminated: English alphabets (A-Z, a-z), Arabic numerals (0-9), mixed-language characters and signs, platform specific items (e.g., emoticons, emojis which are not applicable in the news context).
c)- Punctuation normalization: Systematic removal of both Urdu-specific and generic punctuation marks was done to form word-level tokens with no interference of boundaries: generic punctuation: ‘.‘ (period), ‘,‘ (comma), ‘;‘ (semicolon), ‘:‘ (colon), quotation marks, brackets of any type, Urdu comma, Urdu semicolon, Urdu question mark.
d)- Whitespace standardization: Several successive whitespace characters (spaces, tabs, newlines) were contracted to one, and initial and final whitespace was riddled out. This normalization guarantees that there is a consistency in the token boundaries when further processing is done.
The effect of these cleaning processes is a uniform Urdu text corpus in which semantic information is maintained and structural noise is reduced. This clean text is ideal input to any traditional NLP pipeline as well as modern embedding models.
@@ IV-C4 Urdu stop-word elimination
The text representations can be dominated by stop words that do not provide meaningful information. In the case of Urdu, where the grammatical particles and connectors are abundant, the removal of stop-words is especially effective. A list of 127 Urdu stop words was created manually as a result of linguistic analysis and frequency distribution research of the corpus. A stop-word list contains multiple categories of function words:
@@
Figure 9: Categories of Urdu functional words included in the stop-word list
The removal procedure followed the cleaning of the text but preceded any tokenization or feature extraction. This way, we can recognize stop words in their normal form thereby enhancing accuracy in removal. The effectiveness of Urdu stop-word removal can be seen in the three examples presented in Figure 10. In every example, there is the original (before) text, the text after filtering (after), and the number of characters reduced, and this shows that there is a great deal of compression yet the meaning has not been lost.
@@
Stop-word elimination decreases the size of the corpus by approximately 33.5% quantitatively, indicating the high rate of such functional words within the Urdu news. This automatic shrinkage reduces the computational requirements and can improve the performance of the model by focusing on content carrying terms. Figure 11 shows 127 stop words that were identified and eliminated.
@@ IV-C5 Removal of invalid and redundant entries
The assurance of data quality is also essential to creating trustworthy machine learning models. The pre-processing pipeline removes problematic records based on the following criteria:
a)- Detection and Removal of Null Values: Records that lack essential information in significant columns are deleted. Our targets are the content column (Empty or null text fields,) and the Category column (unlabeled or vague categories) and critical metadata (absent values). This ensures that all training examples are fully featured and labeled.
b)- Identifying Duplicate Content: Duplicate articles (because of republication, syndication, or crawl artifacts) can create bias, as they over-represent part of the content. This move keeps diversity in the datasets and eliminates redundancy.
c)- Minimum length filtering: Text entries that are less than 15 characters do not necessarily have enough semantic meaning. They can be fragments of articles, text of navigation, or placeholders mistakenly captured. This is done to eliminate such entries to retain only substantive content.
@@ IV-C6 Post-Processing Dataset Structure
After the entire pre-processing process, the dataset is organized in the most suitable way to be used in downstream activities. It is encoded in a table format and utilizes UTF-8 encoding, which is compatible with Urdu text processing libraries and frameworks. This is a clean labeled dataset that is used as a starting point to all further operations presented in the subsequent sections.
A combination of these filtering operations results in a clean, uniform dataset. Table II shows the quantitative effects of the pre-processing on the dataset properties.
@@ IV-D Proposed Methodology
The ULTRA framework implements a semantic search-based recommendation system tailored for Urdu news. As formalized in Section IV-A, the core challenge is to retrieve semantically relevant articles from a corpus of over 112k Urdu news items based on user queries of varying lengths and semantic granularity. By ”query granularity,” we refer to the level of detail and specificity in a user’s search input, ranging from concise, keyword-like queries, to elaborate, descriptive sentences. However, in our framework, news articles or full-length documents predominantly function as long queries, targeted at the use case where recommendations are generated for content relevant to the article the user is currently viewing.
Our approach addresses this challenge through three interconnected components: (1) adaptive embedding generation that aligns with query granularity, (2) efficient dimensionality reduction for scalable retrieval, and (3) an optimized dual-path retrieval architecture. The entire methodology is structured to trade semantic accuracy and computational efficiency, to enable real-time recommendations with large-scale deployment.
@@ IV-E Adaptive Embedding Generation
The process of semantic retrieval starts with the creation of high quality vector representations of both queries and documents that represent the semantics of the query and document. We used a pre-trained Urdu transformer model, urduhack/roberta-urdu-small, that generates 768-dimensional contextualized embeddings for input tokens [51]. In order to map these token embeddings into a fixed-dimensional representation, for an entire sequence of text, we use a pooling strategy.
@@ IV-E1 Pooling Strategy Analysis
We tested and analyzed three types of pooling, which are mean pooling, max pooling, and CLS pooling. Mean pooling averages all token embeddings, which is a centroid of semantic information. Max pooling extracts the maximum value across each dimension, potentially emphasizing salient features. CLS pooling utilizes the special classification token’s embedding, which is trained during pre-training to capture sentence-level semantics.
@@ IV-E2 Query-Adaptive Dual-Path Architecture
A key insight driving our methodology is the semantic granularity mismatch between user queries and article representations. Short, focused queries as in Figure 2 typically express a specific topic or intent in few words, making them semantically more aligned with article headlines or titles of documents, which are themselves concise summaries, than with the distributed, detailed content of full articles. To address this, we implemented a query-adaptive routing mechanism based on character length, which dynamically selects the most appropriate semantic space for comparison.
@@
It is important to clarify that in the headline pathway, the system uses headline embeddings only for similarity matching. The actual recommendations presented to the user are the corresponding full articles. This design exploits the fact that headlines serve as succinct summaries of their entire articles, enabling efficient retrieval of relevant content without requiring to compare against lengthy article bodies. We store a headline embedding along with metadata i.e., full article text and category of each article. Once a headline is a match, the whole article can be recommended immediately.
When an article’s content exceeds 512 tokens, we implement a chunking strategy with 50-token overlap. We then take the average of the embeddings of these chunks to generate final embedding of the article. This is the way the entire article is captured.
@@ IV-F Dimensionality Reduction for Scalable Retrieval
At a corpus size of approximately 112,000 articles, it was sluggish to search a 768-dimensional space. We considered four dimensionality reduction methods to enhance scalability alongside preserving the semantic relationships: Principal Component Analysis (PCA), Uniform Manifold Approximation and Projection (UMAP), Linear Discriminant Analysis (LDA), and Autoencoders.
@@ IV-F1 Evaluation Methodology
We evaluated the performance of each method by its similarity to the original space. For each query, we extracted the top-k results from the reduced dimensional database and compared them to the top-k results of the original 768-dimensional database. The overlap score, which is the number of common articles found in both lists, was used to measure how well the reduced version retained the original retrieval rankings.
PCA, a linear approach that preserves the maximum variance, was more effective than the non-linear alternatives across content and headline embeddings. It is deterministic, fast and provided the maximum scores in overlap. Hence we selected it in both of the retrieval paths.
@@ IV-F2 Dimensionality Optimization
After selecting best DR technique, we adjusted the reduced dimension $d^{\prime}$ to trade retrieval accuracy with computational efficiency. In the case of content embeddings (long query pathway), we tried 64, 128, and 256 dimensions with 150 to 350 character length queries. The 128 dimensional setting provided the most favorable tradeoff, it retained a large overlap with the original results.
In the case of headline embeddings (short query pathway), we evaluated 98 queries of 100 characters and discovered that 64 was optimal dimension size that maintained the meaning of short headlines without an apparent drop in retrieval quality. The results of dimension-specific optimization indicated that information density differs between lengthy articles and succinct headlines.
Finally, we stored the pooled and reduced embeddings of optimal dimensions into two separate Chroma vector DBs: one containing content column embeddings, and another containing headline column embeddings.
@@ IV-G System Architecture and Implementation
The ULTRA framework applies the methodological elements in an integrated pipeline illustrated in Figure 4. The system handles user queries through the following steps:
@@ IV-G1 Query Processing and Routing
When a user query is entered into the system as $q$, the system first calculates its character length, which is denoted as $\ell(q)$. Depending on the threshold $\theta=150$ (which again is relative to our recommender system using news corpus and may vary across different recommender systems), the query is routed to the correct pathway. Both paths use the same base transformer model but differ in text source, pooling strategy, and reduced dimension.
@@ IV-G2 Embedding Generation and Retrieval
The system uses two distinct vector databases, each specific to the query type. The pipeline applied to a query is dependent on its character length, denoted by $\ell(q)$, relative to the threshold $\theta=150$, which predetermines the pooling strategy, dimensionality reduction setup, and desired database for similarity search.
In case queries have length of at least theta, i.e., $\ell(q)\geq\theta$ (long queries), the system will use the content pathway. Mean pooling is used to encode the query text into a 768-dimensional embedding $\mathbf{v}_{q}$, and then the principal component analysis (PCA) is used to reduce the pooled embeddings to 128 dimensions. This reduced vector $\mathbf{v}^{\prime}_{q}$ is compared against a pre-computed database of content embeddings, formed by the same mean pooling and PCA reduction procedure, using cosine similarity.
In queries whose length is less than the threshold, i.e., $\ell(q)<\theta$ (short queries), the system uses the headline pathway. The query is CLS pooled to create its initial embedding which is further shrunk to 64 dimensions using PCA. This succinct representation is compared to a separate set of headline embeddings, all computed identically (CLS pooling + PCA to 64D), to identify the most similar headlines in terms of semantics.
Both vector databases are stored in ChromaDB with Hierarchical Navigable Small World (HNSW) indexing cofigured for cosine similarity, which allows efficient approximate nearest-neighbor search at scale.